OpenStudy (anonymous):

when mary emptied her bank she found that she had 84 coins,all nickels and dimes. The value of the coins was $7.15. How many dimes did she have. can i see the process thanks

OpenStudy (radar):

Let x =dimes Let 84-x =nickels

OpenStudy (radar):

10x+5(84-x)=715 10x+420-5x=715 5x=295 x=59 dimes 84-59=25 nickels proof 5.90+1.25=7.15 59+25=84 coins

OpenStudy (radar):

Is there any part you do not understand?

OpenStudy (anonymous):

why is X dimes and not the nickels

OpenStudy (radar):

an arbitrary choice on my part. I could of just as easy chose x to represent the number of nickels. Then 84-x would of been the dimes. However the following equations would of been changed to represent the different values the revised equation would of been: 5x+10(84-x) = 715

OpenStudy (radar):

Comprehendo?

OpenStudy (anonymous):

i understand, thanks alot

OpenStudy (radar):

Great. Remember you are in charge solving equations, but be consistent and logical.

OpenStudy (anonymous):

Yes i will remember